OpenStudy (anonymous):

Which number is a solution of the inequality? HELP x(7 – x) > 8 2 -1 0 8

myininaya (myininaya):

replace the x's with 2's and see if the inequality holds true

myininaya (myininaya):

then do the same until you find one that makes the inequality true

myininaya (myininaya):

like for example enter in 2 2(7-2)>8 is this true is 2(7-2) greater than 8?

OpenStudy (anonymous):

substitute the variable with the number

OpenStudy (anonymous):

i THINK ITS a .... am i right

myininaya (myininaya):

is that because you think 2(7-2)=2(5)=10 is greater than 8 if so yeah it happens to be the first choice we tried

OpenStudy (anonymous):

yes i think thats it!

myininaya (myininaya):

the other choices won't work if we enter in -1 for x we see that -1(7-(-1))=-1(8)=-8 and -8>8 is definitely not true trying x=0 0(7-0)=0 0>8 not true and then so on..